# Relaxed Contrastive Learning for Federated Learning

### v1

@@ 4.2 Representation collapse in FL with SCL
Considering matrix $\mathbf{A}\in\mathbb{R}^{m\times n}$ with its singular values $\{\sigma_{1},...,\sigma_{Q}\}$, where $Q=\min(m,n)$. Let $\mathrm{p}_{k}={\sigma_{k}}/{\sum_{i=1}^{Q}|\sigma_{i}|}$. Then, the effective rank of matrix $\mathbf{A}$ is defined as $\exp({H(\mathrm{p}_{1},...,\mathrm{p}_{Q})})=\exp(-\sum_{k=1}^{Q}\mathrm{p}_{k}\log\mathrm{p}_{k})$, where $H(\cdot)$ is the Shannon entropy.
Figure 2 illustrates the impact of SCL on the effective rank in the CIFAR-100 test set. It supports that SCL diminishes the effective rank when compared to the baseline methods, particularly during the early stage of training, and leads to overall representation collapses. Interestingly, SCL does not exhibit dimensional collapse in the centralized setting [^1] We trained a ResNet-18 model with a single client using the whole CIFAR-100 training set. as in Figure 2, which implies that limited and skewed local training data incurs dimensional collapse in SCL.
These collapsed representations exacerbate the transferability of neural networks across heterogeneous tasks and clients. Previous studies [4, 32, 38] have emphasized the close relationship between feature diversity and transferability, highlighting that representation collapses lead to a loss of crucial information beneficial for downstream tasks. To quantitatively analyze this, we employ the variability collapse index [38], $\text{VCI}=1-\frac{\text{Tr}[\Sigma_{T}^{\dagger}\Sigma_{B}]}{\text{rank}(\Sigma_{B})}$, where $\Sigma_{T}$ and $\Sigma_{B}$ denote the total covariance and between-class covariance matrices for a given feature matrix. It provides a robust measurement of transferability in terms of optimal linear probing loss, where lower values denote better transferability. As observed in Figure 2, SCL exhibits higher VCI values during the early stage, indicating limited transferability even in comparison to the baseline. Given that federated learning can be regarded as a continual fine-tuning process across heterogeneous local tasks, the lack of transferability impedes effective collaborative training, resulting in slow convergence and restrictions on performance enhancements. We will discuss strategies for addressing these challenges in the following subsection.
@@
4.3 Relaxed contrastive learning for FL
@@
$\mathcal{L}_{\text{RCL}}(\mathbf{x}_{i},y_{i};\phi)=-\sum_{\begin{subarray}{c}j\neq i,\\ y_{j}=y_{i}\end{subarray}}\Big{\{}\log\frac{\exp\left({\langle\phi}(\mathbf{x}_{i}),\phi(\mathbf{x}_{j})\rangle/{\tau}\right)}{\sum\limits_{k\neq i}\exp\left({\langle\phi(\mathbf{x}_{i}),\phi(\mathbf{x}_{k})\rangle}/{\tau}\right)}$ (9)
$\hskip 8.53581pt+\beta\cdot\mathbbm{1}_{\left\{\mathbf{x}_{j}\in\mathcal{P}(\mathbf{x}_{i})\right\}}\log\big{(}{\exp\left({\langle\phi(\mathbf{x}_{i}),\phi(\mathbf{x}_{j})\rangle}/{\tau}\right)}\big{)}\Big{\}}$ (10)
where $\mathcal{P}(\mathbf{x})=\left\{{\mathbf{x}^{\prime}}|y_{{\mathbf{x}^{\prime}}}=y_{\mathbf{x}},\langle\phi({\mathbf{x}^{\prime}}),\phi(\mathbf{x})\rangle>\lambda\right\}$ represents a set of intra-class samples with the similarity to the anchor $\mathbf{x}$ higher than the threshold $\lambda$ and $\beta$ is a hyperparameter for the divergence term. The second term of Eq. (10) serves to prevent within-class representation collapses, which also promotes the separation of the examples between different classes. This ultimately enhances overall feature diversity and transferability, which is crucial in the context of federated learning. As illustrated in Figure 3, FedRCL facilitates inter-class separation, mitigates dimensional collapse, and improves the transferability of trained models, resulting in early convergence and significant performance improvement.
@@ 4.4 Multi-level contrastive training
Existing contrastive learning approaches [3, 25, 13] concentrate on aligning the feature representations of the last layer, resulting in predominant model updates in deeper layers while having limited influence on lower-layer parameters. As the parameters are aggregated at all levels of model architectures from heterogeneous clients, this can impede effective model aggregation, particularly in the early layers.
To mitigate this issue, we expand the proposed contrastive learning approach to encompass feature representations in earlier layers. Let $\phi_{l}(\mathbf{x})$ denotes the $l^{\text{th}}$ level feature representation of sample $\mathbf{x}$. Then, we construct $\mathcal{L}_{\text{RCL}}$ by aggregating $\frac{1}{L}\sum_{l=1}^{L}\mathcal{L}_{\text{RCL}}(\mathbf{x},y;\phi_{l})$, where $L$ is the number of layers. For a comprehensive overview of our framework, please refer to Algorithm 1.
4.5 Discussion
FedRCL shares some similarities with existing methods in its use of contrastive loss for local updates, but it has clear differences and advantages over them. While most existing works [19, 23, 5] employ contrastive learning to regulate local training towards the global model for consistent local updates, this constraint often leads to suboptimal solutions as the global model is not fully optimized. Our approach is free from this issue, because FedRCL mitigates local deviations by itself so it does not need to align with the global model explicitly. Moreover, some approaches require additional proxies for contrastive learning such as global prototypes [23] or globally shared data [5], necessitating extra communication overhead and full client participation. In contrast, FedRCL does not require any additional overhead and consistently demonstrates strong performance improvement with an extremely low participation rate. Note that transferring such prototypes or data is even vulnerable to privacy concerns as well.
@@ 5.1 Experimental setup
Our experiments employ a ResNet-18 as the backbone network without pretraining, and replace the batch normalization with the group normalization technique [36] as in [8]. We use the SGD optimizer with a learning rate of 0.1, an exponential decay parameter of 0.998, a weight decay of 0.001, and no momentum, following prior works [1, 39, 15]. We set the number of local training epochs to 5 and adjust the batch size to ensure a total of 10 local update iterations at each local epoch across all experiments. For contrastive learning, we extract the outputs of conv1, conv2_x, conv3_x, conv4_x, and conv5_x layers. We set $\lambda$ to 0.7, $\beta$ to 1, and $\tau$ to 0.05 for all experiments if not specified. All our experiments are implemented using the PyTorch framework [26] and executed on the NVIDIA A5000 GPU. Please refer to the supplementary document for further implementation details.
@@ 5.2 Results
We compare the proposed method, FedRCL, with numerous client-side federated learning baselines [22, 29, 21, 19, 15, 18, 23, 33, 42] on the CIFAR and Tiny-ImageNet datasets. Table 1 demonstrates that our framework outperforms all other existing algorithms by huge margins on all standard benchmarks with various settings. Among the baselines, MOON shows meaningful performance improvement, but its gains are limited under severe data heterogeneity, e.g., $\alpha=0.05$. FedLC employs adaptive label margin to mitigate local deviations, but its effectiveness on performance is constrained. FedDecorr exhibits marginal improvement and, in some settings, even degradation. This suggests that unconditional alleviation of dimensional collapse does not necessarily benefit FL. Compared to these works, our method achieves significant performance improvement across all datasets, regardless of the level of data heterogeneity.
@@ 5.3 Analysis
We validate our framework in more challenging settings with lower client participation rates and the larger number of distributed clients. Table 2 presents the robust performance improvement of FedRCL on three benchmarks, where we set a participation rate of 0.02 and vary the number of clients by $\{100,500\}$. All methods suffer from performance degradation compared to Table 1, due to the reduced client data, increased data disparity, and lower participation rate. Particularly, FedProc suffers from a significant performance drop, because it relies on the global class prototypes aggregated from participating clients at each round, which may not accurately represent every class in this extremely low participation setting. MOON exhibits performance degradation compared to FedAvg in some challenging configurations, partly because it utilizes outdated previous local models due to the sparse participation of local clients. Despite these challenges, FedRCL consistently demonstrates outstanding performance across all datasets.
@@
Table 3 compares the effectiveness of various contrastive learning strategies. FedSCL is an ablative model of our framework, which incorporates a naïve supervised contrastive loss into the FedAvg baseline. While FedSCL improves upon the baseline, its gains are moderate and even undergo degradation in the early stage of training. In contrast, our full framework consistently enhances performance throughout the entire learning process, as also observed in Figure 1. We also employ another variant, denoted as FedCL, which adopts a self-supervised contrastive loss [25], but its benefits are minimal. This is partly because the objective of [25] does not directly align with mitigating the local deviations.
@@
Multi-level training
Table 4 presents the ablative results of our framework using only the last-layer feature outputs on the CIFAR-100 under various levels of data heterogeneity, which verifies that employing early-level representations contributes to achieve additional performance gain.
Figure 4: Ablative results by varying the weight of the divergence penalty ($\beta$), which exhibit stability across a wide range.
Sensitivity on the weight of divergence penalty
We conduct the ablative study on the parameter $\beta$ in Eq. (10) on the CIFAR-100 in non-i.i.d. setting with $\alpha=0.1$. Figure 4 presents that increasing $\beta$ facilitates early convergence and yields consistent performance improvement over a wide range of $\beta$, while excessively high values marginally sacrifice the final performance.
Table 5: Integration of FedRCL into various server-side federated learning approaches under non-i.i.d. setting ($\alpha=0.3$).

### v2

@@ 4.2 Representation collapse in FL with SCL
Consider a matrix $\mathbf{A}\in\mathbb{R}^{m\times n}$ with its singular values $\{\sigma_{1},...,\sigma_{Q}\}$, where $Q=\min(m,n)$, and let $\mathrm{p}_{k}={\sigma_{k}}/{\sum_{i=1}^{Q}|\sigma_{i}|}$. Then, the effective rank of matrix $\mathbf{A}$ is defined as $\exp({H(\mathrm{p}_{1},...,\mathrm{p}_{Q})})=\exp(-\sum_{k=1}^{Q}\mathrm{p}_{k}\log\mathrm{p}_{k})$, where $H(\cdot)$ is the Shannon entropy.
Figure 2 illustrates the impact of SCL on the effective rank in the CIFAR-100 test set. It supports that SCL diminishes the effective rank when compared to the baseline methods, particularly during the early stage of training, and leads to overall representation collapses. Interestingly, SCL does not exhibit dimensional collapse in the centralized setting [^2] We trained a ResNet-18 model with a single client using the whole CIFAR-100 training set. as in Figure 2, which implies that limited and skewed local training data incurs the problem in SCL.
These collapsed representations exacerbate the transferability of neural networks across heterogeneous tasks and clients. Previous studies [4, 32, 38] have emphasized the close relationship between feature diversity and transferability, highlighting that representation collapses of trained models hamper maintaining crucial information beneficial for knowledge transfer to downstream tasks. To quantitatively analyze this, we employ the variability collapse index [38], $\text{VCI}=1-\frac{\text{Tr}[\Sigma_{T}^{\dagger}\Sigma_{B}]}{\text{rank}(\Sigma_{B})}$, where $\Sigma_{T}$ and $\Sigma_{B}$ denote the total covariance and between-class covariance matrices for a given feature matrix. It provides a robust measurement of transferability in terms of optimal linear probing loss, where lower values denote better transferability. As observed in Figure 2, SCL yields higher VCI values at the early stage, indicating low transferability even in comparison to the baselines. Given that federated learning can be regarded as a continual fine-tuning process across heterogeneous local tasks, the lack of transferability impedes collaborative training, resulting in slow convergence and limited performance gain. We will discuss strategies for addressing these challenges in the following subsection.
@@
4.3 Relaxed contrastive loss for FL
@@
$\mathcal{L}_{\text{RCL}}(\mathbf{x}_{i},y_{i};\phi)=\sum_{\begin{subarray}{c}j\neq i,\\ y_{j}=y_{i}\end{subarray}}\Big{\{}-\log\frac{\exp\left({\langle\phi}(\mathbf{x}_{i}),\phi(\mathbf{x}_{j})\rangle/{\tau}\right)}{\sum\limits_{k\neq i}\exp\left({\langle\phi(\mathbf{x}_{i}),\phi(\mathbf{x}_{k})\rangle}/{\tau}\right)}$
$\hskip 8.53581pt+\beta\cdot\log\Big{(}{\sum_{\begin{subarray}{c}{\mathbf{x}_{k}\in\mathcal{P}(\mathbf{x}_{i})}\end{subarray}}\exp\Big{(}{\langle\phi(\mathbf{x}_{i}),\phi(\mathbf{x}_{k})\rangle}/{\tau}\Big{)}}+\exp(1/{\tau})\Big{)}\Big{\}}$ (5)
where $\mathcal{P}(\mathbf{x})=\left\{{\mathbf{x}^{\prime}}|y_{{\mathbf{x}^{\prime}}}=y_{\mathbf{x}},\langle\phi({\mathbf{x}^{\prime}}),\phi(\mathbf{x})\rangle>\lambda\right\}$ represents a set of intra-class samples more similar to the anchor $\mathbf{x}$ than the threshold $\lambda$ and $\beta$ is a hyperparameter for the divergence term. The second term of Eq. (5) serves to prevent within-class representation collapses, which also promotes the separation of the examples between different classes. This ultimately enhances overall feature diversity and transferability, which is crucial in the context of federated learning with non-i.i.d. settings. As illustrated in Figure 3, FedRCL facilitates inter-class separation, mitigates dimensional collapse, and improves the transferability of trained models, resulting in early convergence and significant performance improvement.
@@ 4.4 Multi-level contrastive training
Existing contrastive learning approaches [3, 25, 13] concentrate on aligning the feature representations of the last layer, resulting in predominant model updates in deeper layers while having limited influence on lower-layer parameters. To mitigate this issue, we expand the proposed contrastive learning approach to encompass feature representations in earlier layers. Let $\phi_{l}(\mathbf{x})$ denotes the $l^{\text{th}}$ level feature representation of sample $\mathbf{x}$. Then, we construct $\mathcal{L}_{\text{RCL}}$ by aggregating $\frac{1}{L}\sum_{l=1}^{L}\mathcal{L}_{\text{RCL}}(\mathbf{x},y;\phi_{l})$, where $L$ is the number of layers. The comprehensive algorithm of our framework is presented in Algorithm 1.
@@
4.5 Discussion
FedRCL has something common with existing methods incorporating contrastive loss for local updates, but it has clear differences and advantages over them. While most existing works [19, 23, 5] employ contrastive learning to regulate local training towards the global model for consistent local updates, this constraint often leads to suboptimal solutions as the global model is not fully optimized. Our approach is free from this issue, because FedRCL mitigates local deviations by itself so it does not align with the global model explicitly. Some algorithms require proxies for contrastive learning such as global prototypes [23] or globally shared data [5], which rely on extra communication overhead and full client participation. Note that transferring such prototypes or data is vulnerable and incurs privacy concerns. In contrast, FedRCL does not involve any additional overhead and consistently demonstrates strong performance improvement even with an extremely low participation rate.
@@ 5.1 Experimental setup
We adopt a ResNet-18 as the backbone network, where we replace the batch normalization with the group normalization [36] as suggested in [8]. We trained the model from scratch, using the SGD optimizer with a learning rate of 0.1, an exponential decay parameter of 0.998, a weight decay of 0.001, and no momentum, following prior works [1, 39, 15]. The number of local training epochs is set to 5 and the batch size is adjusted to ensure a total of 10 local iterations at each local epoch throughout all experiments. We apply contrastive learning to conv1, conv2_x, conv3_x, conv4_x, and conv5_x layers. Other hyperparameter settings are as follows for all experiments unless specified otherwise: $\lambda=0.7$, $\beta=1$, and $\tau=0.05$. We used the PyTorch framework [26] for implementation and executed on NVIDIA A5000 GPUs. Please refer to the supplementary document for further details about our implementation.
@@ 5.2 Results
We compare the proposed method, FedRCL, with numerous client-side federated learning baselines [22, 29, 21, 19, 15, 18, 23, 33, 42] on the CIFAR and Tiny-ImageNet datasets. Table 1 demonstrates that our framework outperforms all other existing algorithms by large margins on all datasets and experiment settings. Among the baselines, MOON presents meaningful performance improvement, but its gains are marginal under severe data heterogeneity, e.g., $\alpha=0.05$. FedLC employs adaptive label margin to mitigate local deviations, but its impact on performance is limited. FedDecorr exhibits minor improvement but degraded performance in some settings. This implies that the blind mitigation of dimensional collapse is not necessarily helpful for FL. Compared to other works, our algorithm achieves significant performance improvements in all datasets, regardless of the level of data heterogeneity.
@@ 5.3 Analysis
We validate our framework in more challenging scenarios with lower client participation rates and a larger number of distributed clients. Table 2 presents the robust performance improvement of FedRCL on three benchmarks, where a participation rate is 0.02 and the number of clients is set to one of $\{100,500\}$. All methods suffer from performance degradation, compared with the results in Table 1, due to the reduced client data, increased data disparity, and lower participation rate. Particularly, FedProc experiences a significant performance drop, because it relies on the global class prototypes aggregated from participating clients at each round, which may not be accurate in extremely low participation settings. MOON exhibits performance degradation compared to FedAvg in some challenging configurations, partly because it utilizes outdated previous local models due to the sparse participation of local clients. Despite these challenges, FedRCL consistently demonstrates promising performance on all the tested datasets.
@@
Table 3 compares the effectiveness of various contrastive learning strategies, all employing multi-level contrastive training for fair comparisons. FedSCL is an ablative model of our framework, which incorporates a naïve supervised contrastive loss into the FedAvg baseline. While FedSCL improves upon the baseline in general, its gains are moderate and even negative in the early stage of training. In contrast, our full framework consistently enhances performance throughout the entire learning process, as also observed in Figure 1. We also employ another variant, denoted as FedCL, which adopts a self-supervised contrastive loss [25], but its benefits are not salient. This is partly because the objective in [25] does not directly align with the reduction of local deviations.
@@
Multi-level contrastive learning
Table 4 presents the ablative results of FedRCL on CIFAR-100, where the proposed contrastive learning is applied only to the last-layer feature outputs. The results show that FedRCL benefits from the contrastive learning on intermediate representations.
Figure 4: Ablative results by varying the weight of the divergence penalty ($\beta$), which exhibit stability across its wide range.
Sensitivity of the divergence penalty
We study the impact of $\beta$ in Eq. (5) on the performance of FedRCL under a non-i.i.d. setting with $\alpha=0.1$. Figure 4 illustrates that a large $\beta$ leads to early convergence and the improvements are consistent over a wide range of $\beta$, although its excessively high values marginally degrades the final performances.
Table 5: Integration of FedRCL into various server-side federated learning approaches under a non-i.i.d. setting ($\alpha=0.3$).